# A Novel Approach to Balance Convenience and Nutrition in Meals With Long-Term Group Recommendations and Reasoning on Multimodal Recipes and its Implementation in BEACON

### v1

@@ Introduction
Although it is well known that nutritious foods are essential to a person’s health, the actual adherence to dietary requirements is quite poor across the world. In fact, according to a recent meta-survey (Leme et al. 2021), almost 40% of the population across high and low- and medium-income countries do not adhere to their national food-based dietary guidelines, often prioritizing convenience over nutrition needs. Previous studies have shown that adhering to a provided meal plan instead of a self-selected one reduces the risk for adverse health conditions (Metz et al. 1997). Some people prefer getting food recommendations from their friends or family, and others turn to online recommender systems (Yang et al. 2017) or even Large Language Models (LLMs)(Rostami 2024) as they have become easily available in the form of chatbots, which can be inaccurate, misleading, or not wholly informed. For example, authors in (Papastratis et al. 2024) found that ChatGPT alone is not a reliable tool for meal recommendation when assessing ChatGPT-based recommenders for balanced diets in non-communicable diseases (NCDs) patients.
We seek to help the general population decide on meal choices while nudging them towards healthy choices by leveraging data from online recipes, domain knowledge about meals and how they are configured from foods, and user preferences (Figure 1). In doing so, we recognize the reality that people want to explore a variety of foods, and a long-horizon meal recommender can act as a trusted companion seeking to keep the user well informed even when they deviate from nutrition guidelines.
Our contributions are that we: (1) introduce a novel approach to the meal recommendation problem, accounting for variable meal configurations and flexible time horizons, and propose innovative quantitative metrics to evaluate the framework and benchmark its performance against relevant baselines, (2) adopt the multi-modal R3 framework to convert recipes from two prominent fast food chains into R3 representations, leveraging various LLM-based methodologies to ensure robust and accurate transformation, and (3) present the design of the BEACON meal recommender system, showcasing its potential through a compelling and practical use case to effectively balance both convenience and nutrition.
In the remainder of the paper, we provide relevant background in automated recommendations of personalized meals and then discuss our problem formulation, key solution components including data (recipe representation and format conversion) and meal recommendation, and their evaluation. We then describe a prototype implementation of the solution in the BEACON system along with the supported use cases and conclude with a discussion of practical considerations and avenues for future extensions.
@@
Considering the problem of meal plan recommendation, the representation of the food items is a crucial portion of the overall system. Many works utilize text descriptions (Li and Zaki 2020) of food items or recipes, which limits machine readability. Alternatively, we use a structured form of recipes in R3 format (Pallagani et al. 2022) which covers the content (food description) and also the process of preparing it spanning text and image. This allows a meal recommendation system to present more useful information in an end-end system implementation and reason across modalities as well as with content and preparation processes.
The closest to our work is the SousChef system (Ribeiro et al. 2017), which tackles the problem of food recommendation [^1] They call it meals but do not have the rich meal configurations we support, and instead focus on single items. for improving the health of older adults. SousChef utilizes a two-stage rule-based algorithm to filter away incompatible choices and then recommends food using data about user preferences and item ingredients to recommend multi-item food plans based on nutritional needs; we consider a longer horizon problem setting, learning-based methods, and a multi-modal dataset. Eat This Much (EatThisMuch 2024) is another implementation that considers the problem of meal recommendation and provides users the options to input calories, diet preference(s), and the number of meals they would like. However, this work does not disclose the recommendation algorithm nor does it allow for the customization of meal configurations and time horizons that our work does. Refer to Table 1 for a comparison of our work with the aforementioned two works. Additionally, other systems (Zeevi et al. 2015; Forouzandeh et al. 2024; Min, Jiang, and Jain 2019; Zioutos, Kondylakis, and Stefanidis 2023; Pecune, Callebert, and Marsella 2020) have explored various aspects of food recommendation but often do not incorporate a structured representation that includes both the food content and preparation process.
@@
We define the problem statement by explaining the inputs and outputs of the proposed system (Figure 1). Let ($\mathcal{R}$) be the set of all recipes in the R3 representation. Let ($\mathcal{U}$) be the set that contains all user-provided information, including dietary conditions( Healthy or Diabetic), their ternary food preferences (likesDairy, likesNuts, likesMeat), meal plan format constraints ($\mathcal{C}$). $\mathcal{C}$ includes daily meal names corresponding to the number of meals requested per day (such as Mid-Morning Snack or Dinner), the meal components for each represented as a set $MC\subset\{\text{Beverage},\text{Main},\text{Side},\text{Dessert}\}$, and the length of the meal plan in days.  [^2] In our simulated case study, we configure all recommendations to be in the following format (per day): Breakfast (Main Course, Beverage), Lunch (Main Course, Side, Beverage), and Dinner (Main Course, Side, Dessert, Beverage). The duration for recommendations is configured separately for different simulations(min. one day and max. five days. The output of our system, $\mathcal{MP}$, is a meal plan consisting of meals from ($\mathcal{R}$) informed by ($\mathcal{U}$) in JSON format with a schema defined by $\mathcal{C}$, with meal items outputted by one of three methods, each of which is described more in detail in Solution Components: random selection, sequential selection, and bandit-based selection.
Solution Components and BEACON System Implementation
@@
1-shot prompting 3-shot prompting 5-shot prompting Config $s_{jec}$ $s_{syn}$ $s_{sem}$ $s_{ppl}$ Config $s_{jec}$ $s_{syn}$ $s_{sem}$ $s_{ppl}$ Config $s_{jec}$ $s_{syn}$ $s_{sem}$ $s_{ppl}$ $e_{1}$, $t=0.0$ 437 - 0.666 1.06 $e_{1}$, $t=0.0$ 86 - 0.640 1.06 $e_{1}$, $t=0.0$ 93 - 0.766 1.03 $e_{1}$, $t=0.3$ 59 0.361 0.630 1.24 $e_{1}$, $t=0.3$ 212 - 0.677 1.03 $e_{1}$, $t=0.3$ 690 - 0.516 1.02 $e_{1}$, $t=0.7$ 84 - 0.681 2.32 $e_{1}$, $t=0.7$ 52 - 0.629 15.6 $e_{1}$, $t=0.7$ 86 - 0.667 2.98 $e_{1}$, $t=1.0$ 28 - 0.706 47.6 $e_{1}$, $t=1.0$ 111 - 0.686 49.4 $e_{1}$, $t=1.0$ 54 - 0.662 52.6 $e_{2}$, $t=0.0$ 288 - 0.639 1.09 $e_{2}$, $t=0.0$ 221 - 0.700 1.06 $e_{2}$, $t=0.0$ 1283 - 0.701 1.04 $e_{2}$, $t=0.3$ 387 - 0.639 1.13 $e_{2}$, $t=0.3$ 65 - 0.641 1.26 $e_{2}$, $t=0.3$ 28 - 0.717 1.02 $e_{2}$, $t=0.7$ 41 - 0.688 3.29 $e_{2}$, $t=0.7$ 187 - 0.667 2.27 $e_{2}$, $t=0.7$ 578 - 0.604 1.56 $e_{2}$, $t=1.0$ 133 - 0.689 20.1 $e_{2}$, $t=1.0$ 40 - 0.766 46.5 $e_{2}$, $t=1.0$ 95 - 0.735 48.8
@@
We use three different evaluation criteria to evaluate a recommendation. These criteria include assessing the recurrence of items across meals and within meals (duplicate metric, $md$), assessing if the recommended meals satisfy the preferred components (meal coverage metric, $cs$), and assessing if the recommended meals have ingredients that match user preferences (user-constraint metric, $uc$). For each recommendation, we compute the goodness score $G$ as a weighted sum of the individual scores $md$, $cs$, and $uc$, with weights tailored to user preferences (e.g., some users may prefer having duplicates in their meals, while others may not). We elaborate on the three metrics below.
@@
Duplicate Metric ($dm$) Our duplicate metric examines the occurrence of repeated items meals as a meal item duplicate. The meal item duplicate score, denoted as $dm_{i}$ for a particular meal $m_{i}$, measures the ratio of unique items to total items within the meal. For a recommendation comprising multiple meals $m_{1},...,m_{n}$, we calculate the meal item duplicate score $dm$ as the average of all meal item duplicate scores $dm_{1},...,dm_{n}$.
@@
Meal Coverage Metric ($mc$) Our meal coverage metric evaluates the extent to which a meal recommendation aligns with the user’s desired food roles (Main Course, Side Dish, Dessert, or Beverage). To calculate the coverage score for a recommendation, we analyze the presence of recommended meal items corresponding to their role and alignment with user preferences. If a recommended item matches its assigned role and aligns with user preferences (indicated using a weight of $+1$), it positively contributes to the coverage score. Conversely, misaligned recommended items (e.g., recommending a beverage item like soda as a side dish) would incur a penalty on the coverage score. For each meal, $m_{i}$, we calculate a coverage score $mc_{i}$ by taking the ratio of requested roles fulfilled to the number of requested roles. We calculate our final coverage score, $mc$, as the average of all scores $mc_{1},...,mc_{n}$.
@@
User Constraint Metric ($uc$) In addition to specifying the types of food roles they prefer in each meal, users also provide their ingredient preferences, focusing on three key features: dairy content, meat content, and nuts content. These features were chosen to create a minimally functional system, with plans to extend the list of features in the future. Each feature can have a user preference value of $-1$, 00, or $+1$, representing a negative preference (the user prefers meals without this ingredient), a neutral preference (the user is indifferent), and a positive preference (the user prefers meals with this ingredient), respectively. Our system is designed to be flexible, allowing any number of such features to be added or removed based on user requirements.
@@
We manually annotate each of our fifty-two R3 representations with corresponding feature flags, indicating whether an item contains a particular ingredient (e.g., dairy, meat, or nuts). For each meal $m_{i}$, we calculate a user constraint score $uc_{i}$ by comparing the user’s preference with the meal’s ingredient content. If the user’s preference is negative ($-1$) and the meal contains the ingredient, this counts negatively towards the score. Conversely, if the user’s preference is positive ($+1$) and the meal contains the ingredient, this counts positively towards the score. Neutral preferences (00) do not affect the score regardless of the meal’s content. Additionally, if a meal does not contain an ingredient that the user positively prefers, the overall goodness is not penalized unless the user specifies it through a configurable flag. We calculate the final user constraint score $uc$ as the average of $uc_{1},uc_{2},...,uc_{n}$ across all meals in the recommendation.
@@
M0 serves as our baseline method for meal recommendations. In this approach, meals are generated randomly i.e., the selection of items for each meal does not consider any user preferences, dietary restrictions, or allergen information. An advantage of M0 is that it has the potential to introduce users to a diverse array of food choices. By forming meals without any specific criteria, users might encounter new and varied food items that they may not have otherwise considered, thereby broadening their culinary experiences.
M1 introduces a more structured approach compared to M0. In this method, we use our dataset of recipes and rotate through them to recommend meals. This sequential nature ensures that each recipe in the list is eventually recommended, however, like M0, M1 does not take into account any individual user needs. Its primary advantage over M0 is only the avoidance of repetitive randomness.
@@
Figure 2: Meal preference configuration showing Anthony Gibbson’s selected constraints, including demographic information, dietary restrictions, and meal type selection
In the following sections, we describe a detailed use case of the BEACON system and consider the ethical, legal, and security ramifications of the deployed system. Our proposed solution can be used in a variety of different use cases, including serving as a meal planner for: (a) diabetic individuals looking to be recommended meals that help them manage their condition, (b) individuals from minority communities looking to be recommended culturally relevant meals, (c) busy professionals seeking convenient and healthy meal options, (d) medical professionals seeking to recommend meal plans to their patients.
Use Case
This walk-through highlights the three main interfaces of the system: Meal Preference Configuration, Meal Plan Presentation, and Dietary Insights and Analytics, as shown in Figures 2, 3, and 4 respectively. To illustrate the system’s functionality, we consider a fictional persona: Anthony Gibbson, an African-American individual managing diabetes. Anthony seeks meal plans that align with his dietary needs and cultural food preferences. Through the application, Anthony is guided to configure his profile, receive tailored meal plans, and monitor his dietary patterns with actionable insights.
@@
Anthony begins his journey by configuring his preferences and health constraints using the interactive interface depicted in Figure 2. He enters his personal information, such as height and weight, specifies ‘diabetes’ as his primary health condition, and chooses ‘Soul Food’ from the cuisine preferences drop-down. Since he has no specific restrictions regarding dairy, meat, or nuts, Anthony leaves the corresponding sliders unchanged. Additionally, he opts not to select a predefined diet, such as keto or vegan. Lastly, Anthony also has the option to configure various meal-specific options, such as choosing the meal names and times.
@@
This configuration is processed by the BEACON algorithm, which considers Anthony’s inputs to generate a personalized meal plan that aligns with his health and cultural preferences. By incorporating demographic-specific considerations and evidence-based health constraints, the system ensures that the recommendations cater to his unique requirements. While this walkthrough focuses on Anthony’s specific use case, the system is designed to cater to a diverse range of users with varying health conditions, dietary restrictions, and cultural preferences.
Figure 3: Generated meal plan for Anthony Gibbson, tailored to his preferences and health conditions. The plan includes a breakdown of meal components and plan evaluation scores, ensuring alignment with Anthony’s requirements
Personalized Meal Plan Presentation
Once Anthony submits his preferences, the system presents him with a tailored meal plan via the Meal Plan Presentation Screen, as shown in Figure 3. For example, his breakfast plan includes waffles as the main course, grits as a side dish, and black coffee as a beverage. Each meal component is carefully chosen to ensure a balance of flavor, cultural relevance, and suitability for diabetes management. Anthony can explore various interactive features on this screen:
  • Regenerating individual meals if he desires alternative options.
  • Viewing the goodness score of his recommendation to see how well the plan matches his preferences
  • Saving the plan, marking meals as favorites, or accessing preparation instructions.
  • Choosing to order ingredients directly or customize meals further.
The dashboard provides an intuitive breakdown of meal components, making it easy for Anthony to evaluate and adjust his plan as needed.
Dietary Insights and Analytics
As Anthony continues to use the application, he can monitor his dietary patterns and health metrics through the Analytics Dashboard, shown in Figure 4. This interface offers a personalized view of his progress and dietary impact, including:
  • Daily carbohydrate intake, tracked to ensure it remains within recommended levels for diabetes management.
  • Indicators for hydration levels, meal diversity scores, and blood sugar stability.
  • A visual summary of generated meal plans and their compliance with his health goals.
Interactive tools allow Anthony to explore expandable visualizations, compare the nutritional composition of meals, and identify trends in his dietary habits.
Figure 4: Analytics dashboard visualizing Anthony’s dietary patterns, including daily carbohydrate intake and blood sugar levels, along with providing actionable insights for optimizing his future meal choices
@@
To minimize system abuse, recipe sharing is disabled by default, and all recipe data is vetted for accuracy, and sourced from verified databases such as MyPlate (USDA ). Health disclaimers encourage users to consult professionals for dietary advice, further reducing the risks of misinformation. BEACON complies with all applicable technology licenses, including the MIT and BSD-3-Clause licenses. Health data management adheres to HIPAA standards, and demographic data complies with GDPR (Union 2016) and CCPA (Legislature 2018). Intellectual property is managed under the MIT license and university policies, supported by a Contributor License Agreement (CLA) for transparency.
@@
See the appendix for a detailed description of how we plan to implement the system as a web application.
@@
We conduct our experiments for each recommendation method across 3 different time frames $t_{1}$ (1 day) $t_{2}$ (3 days), and $t_{3}$ (5 days) and 3 user configurations, $c_{1},c_{2},$ and $c_{3}$. These user configurations all include 24 users. For each of the food features that we consider (hasDairy, hasMeat, hasNuts), we consider the corresponding user features $\{u_{f}\}$ that can each take on a value of -1, 0, or 1. These values correspond to negative preference, neutral preference, and positive preference respectively. For each $c_{i}$ and feature $u_{f}$, we randomly select $p_{i}$ users to have a positive preference to $u_{f}$, $n_{i}$ users to have a negative preference to $u_{f}$, and the remaining to have a neutral preference to $u_{f}$. In $c_{1}$, $c_{2}$, and $c_{3}$, we choose $n_{1}=p_{1}=12$, $n_{2}=p_{2}=8$, and $n_{3}=p_{3}=2$. Thus, the constraints on users’ preferences are decreasing across the configurations. Each $c_{i}$ is referred to as $n_{i}$/(24 - $n_{i}$ $p_{i}$)/$p_{i}$ in Figure 5, corresponding to negative, neutral, and positive preference. For each experiment that we conducted, we display our three metrics: user constraint ($uc$), duplicate meal ($dm$), and meal coverage ($mc$), as well as their average ($uc\_dm\_mc$). We display our results graphically in Figure 5. Please refer to Table 5 for an equivalent tabular view.
As shown in Figure 5, M2 expectedly outperforms other methods in the user constraint metric and meal coverage metric as it is the most informed out of the three. It is important to note that when there are fewer users with negative preferences towards the food features, M0 and M1 only perform marginally worse because users are less particular about their preferences. M1 always scores perfectly in the duplicate meal metric because there are more items in our dataset than are in a meal, while M2 performs the worst, which is caused by M2 favoring very few items with a higher probability of being a positively recommended item. This causes M2 to perform poorly in the $uc\_dm$ combination metric for most trials. Additionally, in the $uc\_dm$ metric, M2 performs the worst in the $c_{3}$ configuration because users are less particular and M2 is more likely to recommend duplicates. However, since, M0 and M1 do not perform nearly as well in the $mc$ and $uc$ metrics, M2 performs significantly better in the $uc\_dm\_mc$, $uc\_mc$, and $dm\_mc$ combination metrics. We can also see that M2 is a precise and accurate method regardless of the user configuration and M0 and M1 are somewhat accurate for some metrics, but lack precision. We thus conclude that the boosted bandit algorithm represented by M2 is the most effective in recommending meals out of the three.

### v2

@@ Introduction
Although it is well known that nutritious foods are essential to a person’s health, the actual adherence to dietary requirements is quite poor across the world. In fact, according to a recent meta-survey (Leme et al. 2021), almost 40% of the population across high-, low-, and medium-income countries do not adhere to their national food-based dietary guidelines, often prioritizing convenience over nutrition needs. Previous studies have shown that adhering to a provided meal plan instead of a self-selected one reduces the risk for adverse health conditions (Metz et al. 1997). Some people prefer getting food recommendations from their friends or family, and others turn to online recommender systems (Yang et al. 2017) or even Large Language Models (LLMs) (Rostami 2024) as they have become easily available in the form of chatbots, albeit being inaccurate, misleading, or not wholly informed. For example, authors in (Papastratis et al. 2024) found that ChatGPT alone is not a reliable tool for meal recommendation when assessing ChatGPT-based recommenders for balanced diets in non-communicable disease (NCD) patients.
We seek to help the general population decide on meal choices while nudging them towards healthy choices by leveraging data from online recipes, domain knowledge about meals and how they are configured from foods, and user preferences (Figure 1). In doing so, we recognize the reality that people want to explore a variety of foods, and a long-horizon meal recommender can act as a trusted companion seeking to keep the user well-informed even when they deviate from nutrition guidelines.
Our contributions are that we: (1) introduce a novel approach to the meal recommendation problem, accounting for variable meal configurations and flexible time horizons, and propose innovative quantitative metrics to evaluate the framework and benchmark its performance against relevant baselines, (2) adopt the multi-modal R3 framework (Pallagani et al. 2022) to convert recipes from two prominent fast food chains into R3 representations, leveraging various LLM-based methodologies to ensure robust and accurate transformation, and (3) present the design of the BEACON meal recommender system, showcasing its potential through a compelling and practical use case to effectively balance both convenience and nutrition.
In the remainder of the paper, we provide relevant background in automated recommendations of personalized meals and then discuss our problem formulation, key solution components including data (recipe representation and format conversion) and meal recommendation, and their evaluation. We then describe a prototype implementation of the solution in the BEACON system, along with the supported use cases, and conclude with a discussion of practical considerations and avenues for future extensions.
@@
Considering the problem of meal plan recommendation, the representation of the food items is a crucial portion of the overall system. Many works utilize text descriptions (Li and Zaki 2020) of food items or recipes, which limits machine readability. Alternatively, we use a structured form of recipes in R3 format (Pallagani et al. 2022) which covers the content (food description) and also the process of preparing it, spanning text and image. This allows a meal recommendation system to present more useful information in an end-to-end system implementation and reason across modalities as well as with content and preparation processes.
The closest to our work is the SousChef system (Ribeiro et al. 2017), which tackles the problem of food recommendation [^1] They call it meals but do not have the rich meal configurations we support, and instead focus on single items. for improving the health of older adults. SousChef utilizes a two-stage rule-based algorithm to filter away incompatible choices and then uses user preference data and item ingredients to recommend multi-item food plans based on nutritional needs; we consider a longer horizon problem setting, learning-based methods, and a multi-modal dataset. Eat This Much (EatThisMuch 2024) is another implementation that considers the problem of meal recommendation and provides users with the options to input calories, diet preference(s), and the number of meals they would like. However, this work does not disclose the recommendation algorithm, nor does it allow for the customization of meal configurations and time horizons that our work does. Refer to Table 1 for a comparison of our work with the aforementioned two works. Additionally, other systems (Zeevi et al. 2015; Forouzandeh et al. 2024; Min, Jiang, and Jain 2019; Zioutos, Kondylakis, and Stefanidis 2023; Pecune, Callebert, and Marsella 2020) have explored various aspects of food recommendation, but often do not incorporate a structured representation that includes both the food content and preparation process.
@@
We define the problem statement by explaining the inputs and outputs of the proposed system (Figure 1). Let $\mathcal{R}$ be the set of all recipes in the R3 representation. Let $\mathcal{U}$ be the set that contains all user-provided information, including dietary conditions (Healthy or Diabetic), their ternary food preferences (likesDairy, likesNuts, likesMeat), and meal plan format constraints $\mathcal{C}$. $\mathcal{C}$ includes daily meal names corresponding to the number of meals requested per day (such as Mid-Morning Snack or Dinner), the meal components for each represented as a set $MC\subset\{\text{Beverage},\text{Main},\text{Side},\text{Dessert}\}$, and the length of the meal plan in days.  [^2] In our simulated case study, we configure all recommendations to be in the following format (per day): Breakfast (Main Course, Beverage), Lunch (Main Course, Side, Beverage), and Dinner (Main Course, Side, Dessert, Beverage). The duration for recommendations is configured separately for different simulations (min one day and max five days). The output of our system $\mathcal{MP}$ is a meal plan consisting of meals from $\mathcal{R}$ informed by $\mathcal{U}$ in JSON format with a schema defined by $\mathcal{C}$, with meal items outputted by one of three methods, detailed in Solution Components section: random selection, sequential selection, and bandit-based selection.
Solution Components and
BEACON System Implementation
@@
1-shot prompting 3-shot prompting 5-shot prompting Config $_{jec} ss$$_{syn}$ $s_{sem}$ $s_{ppl}$ Config $_{jec} ss$$_{syn}$ $s_{sem}$ $s_{ppl}$ Config $_{jec} ss$$_{syn}$ $s_{sem}$ $s_{ppl}$ $e_{1}$, $t=0.0$ 437 - 0.666 1.06 $e_{1}$, $t=0.0$ 86 - 0.640 1.06 $e_{1}$, $t=0.0$ 93 - 0.766 1.03 $e_{1}$, $t=0.3$ 59 0.361 0.630 1.24 $e_{1}$, $t=0.3$ 212 - 0.677 1.03 $e_{1}$, $t=0.3$ 690 - 0.516 1.02 $e_{1}$, $t=0.7$ 84 - 0.681 2.32 $e_{1}$, $t=0.7$ 52 - 0.629 15.6 $e_{1}$, $t=0.7$ 86 - 0.667 2.98 $e_{1}$, $t=1.0$ 28 - 0.706 47.6 $e_{1}$, $t=1.0$ 111 - 0.686 49.4 $e_{1}$, $t=1.0$ 54 - 0.662 52.6 $e_{2}$, $t=0.0$ 288 - 0.639 1.09 $e_{2}$, $t=0.0$ 221 - 0.700 1.06 $e_{2}$, $t=0.0$ 1283 - 0.701 1.04 $e_{2}$, $t=0.3$ 387 - 0.639 1.13 $e_{2}$, $t=0.3$ 65 - 0.641 1.26 $e_{2}$, $t=0.3$ 28 - 0.717 1.02 $e_{2}$, $t=0.7$ 41 - 0.688 3.29 $e_{2}$, $t=0.7$ 187 - 0.667 2.27 $e_{2}$, $t=0.7$ 578 - 0.604 1.56 $e_{2}$, $t=1.0$ 133 - 0.689 20.1 $e_{2}$, $t=1.0$ 40 - 0.766 46.5 $e_{2}$, $t=1.0$ 95 - 0.735 48.8
@@
We use three different evaluation criteria to evaluate a recommendation. These criteria include assessing the recurrence of items across meals and within meals (duplicate metric, $md$), assessing if the recommended meals satisfy the preferred components (meal coverage metric, $cs$), and assessing if the recommended meals have ingredients that match user preferences (user-constraint metric, $uc$). For each recommendation, we compute the goodness score $G$ as a weighted sum of the individual scores $md$, cscs, and $uc$, with weights tailored to user preferences (e.g., some users may prefer having duplicates in their meals, while others may not). We elaborate on the three metrics below.
@@
Duplicate Metric ($dm$) Our duplicate metric examines the occurrence of repeated items in meals as a meal item duplicate. The meal item duplicate score, denoted as $dm_{i}$ for a particular meal $m_{i}$, measures the ratio of unique items to total items within the meal. For a recommendation comprising multiple meals $m_{1},...,m_{n}$, we calculate the meal item duplicate score $dm$ as the average of all meal item duplicate scores $dm_{1},...,dm_{n}$.
@@
Meal Coverage Metric ($mc$) Our meal coverage metric evaluates the extent to which a meal recommendation aligns with the user’s desired food roles (Main Course, Side Dish, Dessert, or Beverage). To calculate the coverage score for a recommendation, we analyze the presence of recommended meal items corresponding to their role and alignment with user preferences. If a recommended item matches its assigned role and aligns with user preferences (indicated using a weight of +1+1), it positively contributes to the coverage score. Conversely, misaligned recommended items (e.g., recommending a beverage item like soda as a side dish) would incur a penalty on the coverage score. For each meal, $m_{i}$, we calculate a coverage score $mc_{i}$ by taking the ratio of requested roles fulfilled to the number of requested roles. We calculate our final coverage score, $mc$, as the average of all scores $mc_{1},...,mc_{n}$.
@@
User Constraint Metric ($uc$) In addition to specifying the types of food roles they prefer in each meal, users also provide their ingredient preferences, focusing on three key features: dairy content, meat content, and nut content. These features were chosen to create a minimally functional system, with plans to extend the list of features in the future. Each feature can have a user preference value of 1-1, 0, or +1+1, representing a negative preference (the user prefers meals without this ingredient), a neutral preference (the user is indifferent), and a positive preference (the user prefers meals with this ingredient), respectively. Our system is designed to be flexible, allowing any number of such features to be added or removed based on user requirements.
@@
We manually annotate each of our fifty-two R3 representations with corresponding feature flags, indicating whether an item contains a particular ingredient (e.g., dairy, meat, or nuts). For each meal $m_{i}$, we calculate a user constraint score $uc_{i}$ by comparing the user’s preference with the meal’s ingredient content. If the user’s preference is negative (1-1) and the meal contains the ingredient, this counts negatively towards the score. Conversely, if the user’s preference is positive (+1+1) and the meal contains the ingredient, this counts positively towards the score. Neutral preferences (0) do not affect the score regardless of the meal’s content. Additionally, if a meal does not contain an ingredient that the user positively prefers, the overall goodness is not penalized unless the user specifies it through a configurable flag. We calculate the final user constraint score $uc$ as the average of $uc_{1},uc_{2},...,uc_{n}$ across all meals in the recommendation.
@@
M0 serves as our baseline method for meal recommendations. In this approach, meals are generated randomly, i.e., the selection of items for each meal does not consider any user preferences, dietary restrictions, or allergen information. An advantage of M0 is that it has the potential to introduce users to a diverse array of food choices. By forming meals without any specific criteria, users might encounter new and varied food items that they may not have otherwise considered, thereby broadening their culinary experiences.
M1 introduces a more structured approach compared to M0. In this method, we use our dataset of recipes and rotate through them to recommend meals. This sequential nature ensures that each recipe in the list is eventually recommended; however, like M0, M1 does not take into account any individual user needs. Its primary advantage over M0 is only the avoidance of repetitive randomness.
@@
Figure 2: Left: Meal preference configuration showing Anthony Gibbson’s selected constraints, including demographic information, dietary restrictions, nutritional goals, and meal type selection. Right: The interactive, calendar-based meal visualization for Anthony Gibbson, tailored to his preferences and health conditions. The system allows Anthony to build, visualize, realize, and track his meal plans. Arrows and annotated text are presented for clearer explanations.
In the following sections, we describe a detailed use case of the BEACON system and consider the ethical, legal, and security ramifications of the deployed system. Our proposed solution can be used in a variety of different use cases, including serving as a meal planner for (a) diabetic individuals looking to be recommended meals that help them manage their condition, (b) individuals from minority communities looking to be recommended culturally relevant meals, (c) busy professionals seeking convenient and healthy meal options, and (d) medical professionals seeking to recommend meal plans to their patients.
Use Case: Managing Diabetes
This walk-through highlights the two main interfaces of the system: Meal Preference Configuration and Meal Calendar Presentation, as shown in Figure 2. To illustrate the system’s functionality, we consider a fictional persona: Anthony Gibbson, an African-American individual managing diabetes. Anthony seeks meal plans that align with his dietary needs and cultural food preferences. Through our application, Anthony is guided to configure his profile, receive tailored meal plans, and monitor his dietary patterns with actionable insights. Detailed figures of the system are given in the appendix.
@@
Anthony begins his journey on the Food Preferences page, the interactive interface depicted in Figure 2 left. Here, he has the option to select a predefined persona that populates the page with default values or, as in this walk-through, to enter his information manually for a fully personalized experience.
He populates the form with his anthropometric data (such as height and weight), dietary preferences, and specific nutritional goals. For his dietary preferences, he specifies ‘diabetes’ as his primary health condition and, since he has no specific restrictions regarding dairy, meat, or nuts, Anthony leaves the corresponding sliders unchanged. Additionally, he opts not to select a predefined persona, such as a college student or a worklift operator. To further tailor the plan, he sets his nutritional goals, like limiting daily carbohydrate intake, and configures meal-specific options, such as meal names and times.
@@
The BEACON algorithm processes these inputs, incorporating demographic and health constraints to generate a personalized meal plan. While this example focuses on Anthony, the system is designed for a diverse range of users with varying health conditions and cultural preferences.
Personalized Meal Calendar Presentation
Once Anthony submits his preferences, the system presents him with a tailored meal plan via the interactive Meal Calendar, as shown in Figure 2 right. This interface moves beyond static lists, offering a dynamic environment where Anthony can build, visualize, realize, and track his meal plans over time. This interface is designed with multiple views to give him full control over how he visualizes and interacts with his plan.
By default, he is presented with the Meal View, which shows his recommended meals sequentially. From this view, he can accept a recommendation to save it to his meal history, reject it, favorite it to influence future suggestions, or regenerate all recommendations for a fresh set of options. Clicking on any meal reveals its key nutritional information, such as calories, carbohydrates, protein, and fiber. For more details, Anthony can switch to the Food Item View. This breaks down each meal into its individual components: main courses, side dishes, and beverages. This is particularly useful if he decides to adhere to only part of a recommendation. For instance, he can click on ‘waffles’ to view its specific nutritional data and access the recipe needed to prepare it. Finally, the Ingredient View provides the most granular level of detail by breaking down each food item into its constituent ingredients. This view serves as a practical tool, giving Anthony a quick glance at all the ingredients he will need to purchase to prepare his scheduled meals for the week.
The system also provides actionable insights and nudges to help him adhere to his goals for diabetic management. For instance, if it detects that he schedules high-sugar meals for several consecutive days, it might visually flag those entries and suggest healthier, culturally relevant alternatives. This behavioral feedback loop learns from his choices to refine future recommendations, ensuring they are both effective and sustainable.
These views empower Anthony to interact with his plan in several ways: (i) Build and Replan: He can dynamically adjust the plan by manually editing meals, accepting AI-driven suggestions, or regenerating specific meals of entire days if his needs change. (ii) Visualize and Track Goals: The calendar displays his progress towards nutritional targets (e.g., daily carbohydrate limits). Visual cues, such as progress bars or alerts, highlight how each meal impacts his goals. (iii) Realize the Meal Plan: For any given meal, he can choose how to execute it. He can access cooking instructions or opt to order the necessary meal through integrated services. (iv) Analyze Historical Patterns: The system allows him to review past consumption data, helping him recognize recurring habits, such as frequently skipping lunch or relying on the same side dishes, and make informed adjustments.
This intuitive dashboard transforms the meal plan into a comprehensive dietary partner, empowering Anthony to evaluate, adjust, and realize his health objectives.
@@
To minimize system abuse, recipe sharing is disabled by default, and all recipe data is vetted for accuracy and sourced from verified databases such as MyPlate (of Agriculture ). Health disclaimers encourage users to consult professionals for dietary advice, further reducing the risks of misinformation. BEACON complies with all applicable technology licenses, including the MIT and BSD-3-Clause licenses. Health data management adheres to HIPAA standards, and demographic data complies with GDPR (Union 2016) and CCPA (Legislature 2018). Intellectual property is managed under the MIT license and university policies, supported by a Contributor License Agreement (CLA) for transparency.
@@
Refer to the appendix for a detailed description of how we plan to implement the system as a web application.
@@
[…]
As shown in Figure 3, M2 expectedly outperforms other methods in the user constraint metric and meal coverage metric as it is the most informed out of the three. It is important to note that when there are fewer users with negative preferences towards the food features, M0 and M1 only perform marginally worse because users are less particular about their preferences. M1 always scores perfectly in the duplicate meal metric because there are more items in our dataset than are in a meal, while M2 performs the worst, which is caused by M2 favoring very few items with a higher probability of being a positively recommended item. This causes M2 to perform poorly in the uc_dmuc\_dm combination metric for most trials. Additionally, in the uc_dmuc\_dm metric, M2 performs the worst in the $c_{3}$ configuration because users are less particular and M2 is more likely to recommend duplicates. However, since, M0 and M1 do not perform nearly as well in the $mc$ and $uc$ metrics, M2 performs significantly better in the uc_dm_mcuc\_dm\_mc, uc_mcuc\_mc, and dm_mcdm\_mc combination metrics. We can also see that M2 is a precise and accurate method regardless of the user configuration and M0 and M1 are somewhat accurate for some metrics, but lack precision. We thus conclude that the boosted bandit algorithm represented by M2 is the most effective in recommending meals out of the three.